# On Joint Noise Scaling in Differentially Private Federated Learning with Multiple Local Steps

### v1

@@ Abstract
Federated learning is a distributed learning setting where the main aim is to train machine learning models without having to share raw data but only what is required for learning. To guarantee training data privacy and high-utility models, differential privacy and secure aggregation techniques are often combined with federated learning. However, with fine-grained protection granularities the currently existing techniques require the parties to communicate for each local optimisation step, if they want to fully benefit from the secure aggregation in terms of the resulting formal privacy guarantees. In this paper, we show how a simple new analysis allows the parties to perform multiple local optimisation steps while still benefiting from joint noise scaling when using secure aggregation. We show that our analysis enables higher utility models with guaranteed privacy protection under limited number of communication rounds.
@@ 1 Introduction
Federated learning (FL) [31, 25] is a common distributed learning setting, where a central server and several clients holding their own local data sets collaborate to learn a single global model. The main feature in FL is that the clients do not directly communicate data, but only what is required for learning, e.g., gradients or updated model parameters (pseudo-gradients).
@@
While FL satisfies the data minimization principle, i.e., only what is actually needed is communicated while the actual raw data never leaves the client, it does not protect against privacy attacks such as membership inference [38] or reconstruction [19, 51]. Instead, training data privacy is commonly ensured by combining differential privacy (DP) [17]), a formal privacy definition, and secure multiparty computation (MPC), [50] with FL (see, e.g., [25]).
@@
DP is essentially a robustness guarantee for stochastic algorithms, which guarantees that small perturbations to the inputs have small effects on the algorithms’ output probabilities. What constitutes a small perturbation depends on the chosen protection granularity: the same basic DP definition can be used for ensuring privacy on anything from single sample to entire data set level. In turn, MPC protocols can be used to limit the amount of information an adversary has about computations. In FL, secure aggregation (SecAgg) protocols, a specialised form of secure computation that requires significantly less resources than general MPC, are commonly used for communicating model updates from the clients to the server, which can result in provably better joint DP guarantees than is possible to achieve by any single client in isolation.
@@
Under the general FL setup, two main alternatives are commonly considered: cross-device FL and cross-silo FL. In cross-device FL, each client is assumed to have a small local data set, while the total number of clients is large, e.g., thousands or millions. In the cross-silo case, the total number of clients is small, for example, a dozen, but each client is assumed to have a larger local data set. In this paper, our running example is standard cross-silo differentially private FL (DPFL) where the clients communicate all updates to the server using SecAgg. [^1] Instead of considering any specific SecAgg implementation, in this work we mostly assume an idealised trusted aggregator. In this setting, the most useful DP protection granularity is typically something strictly more fine-grained than client-level: when clients are, e.g., different hospitals or banks, there are typically several individuals in a single clients’ local data set and the protection granularity needs to match the use case.
@@
While client-level granularity in DPFL is, at least in principle, straightforward to combine with SecAgg, more fine-grained granularities present a problem: using existing techniques one has to choose between i) having joint DP guarantees with less noise due to SecAgg but with all clients using only a single local optimisation step per FL round, and ii) having more noisy local DP (LDP) guarantees that do not formally benefit from SecAgg while allowing the clients to do more local optimisation steps per FL round. Both of these options have significant drawbacks: the amount of server-client communications is typically one of the first bottlenecks that limit model training in FL, while LDP guarantees regularly require noise levels that heavily affect the resulting model utility. In this paper we show that this trade-off is not unavoidable but can be largely remedied by a simple new analysis of the problem.
@@
  • We present a novel analysis showing when we can increase the number of local optimisation steps in FL using fine-grained DP granularity, while still benefiting from joint DP guarantees using a trusted aggregator.
  • We demonstrate empirically that the proposed approach can lead to large utility benefits without requiring any changes to the underlying algorithms under various settings.
@@ 2 Related Work
There is a significant amount of existing work focusing on the general problem of combining DP with FL, although the focus has mostly been on the cross-device FL setting with user-level DP. To the best of our knowledge, while the combination of DPFL with SecAgg is certainly not novel (see, e.g., [45, 25, 21, 42, 49]), there is no existing work on the privacy analysis when the clients do multiple local optimisation steps fine-grained sample-level DP granularity and communicate the results via SecAgg. [^2] Note that [45, Algorithm 4] seem to state a weaker, specialised version of our results, i.e., they use several local optimisation steps with sample-level DP and SecAgg in FL, while scaling the noise jointly over the clients. However, as noted in [30], the approach in [45] would require a separate proof of privacy beyond what is actually provided in the paper.
@@
Considering the existing work in more detail, we can distinguish several main lines of research. There are many papers proposing novel learning methods for FL, assuming sample-level DP and joint noise scaling with SecAgg. While the existing work only uses a single local optimisation step (see, e.g., [21, 30, 42, 49]), our analysis can be utilised in this setting to enable multiple local steps generally for many such methods.
@@
Another clear line of work has focused on introducing novel discrete DP mechanisms that can be used with additively homomorphic encryption techniques, which typically operate on the group of integers with modulo additions. [2] proposed a binomial mechanism that provides DP using discrete binomial noise. Improving on the binomial mechanism, [11] proposed a discrete Gaussian mechanism, while [1] introduced a Skellam mechanism and [14] a Poisson-binomial mechanism, both of which improve on the discrete Gaussian, e.g., by being infinitely divisible distributions: the sum of Skellam/Poisson-binomial distributed random variables is another Skellam/Poisson-binomial random variable. Our work is not focused on introducing new DP mechanisms, but our analysis allows for using many different DP noise mechanism. In particular, our analysis allows for joint noise scaling under SecAgg including when using infinitely divisible DP mechanisms, such as the Skellam mechanism, with pseudo-gradients and fine-grained DP protection level.
One theme that has seen much recent activity is the combination of DPFL with quantization techniques to minimize the amount of bits communicated by the clients. [2] combined quantisation with the binomial mechanism, while [24] used discrete Gaussian with randomised rounding for DPFL, and [1] did the same with the Skellam mechanism. Again, one main motivation stated for introducing the Poisson-binomial mechanism [14] is that it requires the clients to communicate less bits as the privacy gets tighter, contrary to what happens, e.g., with the discrete Gaussian mechanism. Furthermore, in the tight communication budget regime, [23] proposed stochastic DP sign-SGD, which provides LDP guarantees and communicates 1 bit per client gradient dimension. In the same vein, [12] proposed the Minimum Variance Unbiased (MVU) mechanism for combining DPFL and heavy compression, while [20] introduced Interpolated MVU (I-MVU), that improves on MVU for DPFL especially in the case of high-dimensional data by relaxing the requirement of unbiasedness. While in principle it is possible to combine our analysis with quantisation techniques which are compatible with SecAgg, since the model updates in our case are pseudo-gradients and not gradients, this is not entirely straightforward.
Another alternative approach to improve communication-efficiency of DPFL is to compress the updates directly instead of or in addition to quantising them. On this front, [44] introduced an approach based on relative entropy coding, where the DP guarantees are derived based on the inherent randomness of the compression method without needing any additional DP noise mechanism. However, their approach does not support using SecAgg for joint noise scaling but results in LDP guarantees, and so is not directly compatible with our approach. Finally, [13] proposed and analysed a linear compression scheme, that is compatible with SecAgg, based on sparse random projections. Due to linearity, our approach should be compatible with their method, and lead to reduced communications while still allowing for joint noise scaling.
We leave the detailed consideration and comparison of the possible methods for reducing the required communication budget beyond what is possible by pushing more optimisation steps to the clients while reducing FL rounds for future work.
@@
In summary, while many of the contributions cited above, e.g., novel DP mechanisms, are not limited to cross-device FL, all the experiments and use cases mentioned in the cited papers that are compatible with SecAgg and use multiple local steps only consider joint noise scaling with user- or client level DP in cross-device FL. In contrast, we focus on more fine-grained DP granularities, e.g., on sample-level DP. As we discuss in Section 3, combining sample-level DP with multiple local steps and joint noise scaling using SecAgg with good utility requires a novel privacy analysis. The main aim of this paper is to provide such an analysis.
@@ 3 Background
Federated learning [31, 25] is a collaborative learning setting, where the participants include a central server and clients holding some data. On each FL round, the server chooses a group of clients for an update and sends them the current model parameters. The chosen clients update their local model parameters by taking some amount of optimisation steps using only their own local data, and then send an update back to the server. The server then aggregates the client-specific contributions to update the global model. We use the standard federated averaging update rule: assuming w.l.o.g. that clients $i=1,\dots,N$ have been selected at FL round $t$, and that client $i$ sends an update $\Delta^{(t)}_{i}$ (pseudo-gradient), the updated global model $\theta_{t}$ is given by
@@
We want to guarantee privacy of the trained model w.r.t. the training data, for which we use DP. Writing the space of possible data sets as $\mathcal{X}^{*}:=\cup_{n\in\mathbb{N}}\mathcal{X}^{n}$, we have the following:
@@
[17, 16] Let $\varepsilon>0$ and $\delta\in[0,1]$. A randomised algorithm ${\mathcal{A}}:\,\mathcal{X}^{*}\rightarrow\mathcal{O}$ is $(\varepsilon,\delta)$-DP if for every $x,x^{\prime}\in\mathcal{X}^{*}:x\simeq x^{\prime}$, and every measurable set $E\subset\mathcal{O}$,
@@
$H_{\alpha}(P\|Q):=$ $\mathbb{E}_{t\sim Q}\left(\left[\frac{dP}{dQ}(t)-\alpha\right]_{+}\right)$ (2)
$=$ $\mathbb{E}_{t\sim Q}\left(\left[\frac{p(t)}{q(t)}-\alpha\right]_{+}\right),$ (3)
@@
It has been shown that a randomised algorithm ${\mathcal{A}}$ is $(\varepsilon,\delta)$-DP iff $\sup_{x\simeq x^{\prime}}H_{{\rm e}\hskip 1.0pt^{\varepsilon}}({\mathcal{A}}(x)\|{\mathcal{A}}(x^{\prime}))\leq\delta$ [5, 6].
Our main results do not depend on the exact neighbourhood definition, but in all the experiments we use the add/remove relation or unbounded DP, that is, $x,x^{\prime}\in\mathcal{X}^{*}$ are neighbours, if $x$ can be transformed into $x^{\prime}$ by adding or removing a single protected unit from $\mathcal{X}$. For the protection granularity, although this is not a strict limitation, in the experiments we focus on sample-level DP, i.e., a single protected unit corresponds to a single sample of data. As noted earlier, our analysis could be advantageous for anything more fine-grained than client-level DP, e.g., element-level DP [4] or even for individual-level when there are more than one individual in the clients’ local data.
@@
[53] A pair of distributions $(P,Q)$ is a dominating pair for a stochastic algorithm ${\mathcal{A}}$, if for all $\alpha\geq 0$
@@
For privacy accounting we also utilize Rényi DP (RDP):
Definition 3.4.
[33] Let $\alpha>1$ and $\varepsilon>0$. A randomised algorithm ${\mathcal{A}}:\,\mathcal{X}^{*}\rightarrow\mathcal{O}$ is $(\alpha,\varepsilon)$-RDP if for every $x,x^{\prime}\in\mathcal{X}^{*}:x\simeq x^{\prime}$
$D_{\alpha}({\mathcal{A}}(x)\|{\mathcal{A}}(x^{\prime}))\leq\varepsilon,$
where $D_{\alpha}$ is the Rényi divergence of order $\alpha$:
For plotting all experimental results, we convert RDP into ADP (see Appendix Appendix for more details on privacy accounting).
While DP offers strict privacy protection, it comes at the cost of reduced model utility. This is especially true in the LDP setting, where each client protects their own data independently of any other party [26]. One well-known technique to improve model utility in DPFL has been to utilise SecAgg to turn LDP guarantees into joint guarantees or distributed DP guarantees that depend on multiple clients (see, e.g., [25]). However, naively combining fine-grained DP protection with SecAgg for distributed DP runs into problems, as we demonstrate in the rest of this section.
Starting with the unproblematic case of client-level DP, writing $TA$ for an ideal trusted aggregator and using the well-known Gaussian mechanism [16], one can get joint DP guarantees for any number of local optimisation steps with the following update:

### v2

# On Using Secure Aggregation in Differentially Private Federated Learning with Multiple Local Steps

@@ Abstract
Federated learning is a distributed learning setting where the main aim is to train machine learning models without having to share raw data but only what is required for learning. To guarantee training data privacy and high-utility models, differential privacy and secure aggregation techniques are often combined with federated learning. However, with fine-grained protection granularities, e.g., with the common sample-level protection, the currently existing techniques generally require the parties to communicate for each local optimization step, if they want to fully benefit from the secure aggregation in terms of the resulting formal privacy guarantees. In this paper, we show how a simple new analysis allows the parties to perform multiple local optimization steps while still benefiting from using secure aggregation. We show that our analysis enables higher utility models with guaranteed privacy protection under limited number of communication rounds.
@@ 1 Introduction
Federated learning (FL; McMahan et al. 2017; Kairouz et al. 2019) is a common distributed learning setting, where a central server and several clients holding their own local data sets collaborate to train a single global model. The main feature in FL is that the clients do not directly communicate data, but only what is required for learning, e.g., gradients or updated model parameters (pseudo-gradients).
@@
While FL satisfies the data minimization principle, i.e., only what is actually needed is communicated while the actual raw data never leaves the client, it does not protect against privacy attacks such as membership inference (Shokri et al., 2017) or reconstruction (Fredrikson et al., 2014; Yeom et al., 2018). Instead, training data privacy is commonly ensured by combining differential privacy (DP; Dwork et al. 2006b), a formal privacy definition, and secure multiparty computation (MPC; Yao 1982) with FL (see, e.g., Kairouz et al. 2019).
@@
DP is essentially a robustness guarantee for stochastic algorithms, which guarantees that small perturbations to the inputs have small effects on the algorithms’ output probabilities. What constitutes a small perturbation depends on the chosen protection granularity: the same basic DP definition can be used for ensuring privacy on anything from single sample to entire data set level. In turn, MPC protocols can be used to limit the amount of information an adversary has about computations. In FL, secure aggregation (SecAgg) protocols, a specialised form of secure computation that requires significantly less resources than general MPC, are commonly used for communicating model updates from the clients to the server. This approach enables provably better distributed DP (DDP) guarantees, that rely jointly on a group of clients, than is possible to achieve by any single client in isolation.
@@
Under the general FL setup, two main alternatives are commonly considered: cross-device FL and cross-silo FL (Kairouz et al., 2019). In cross-device FL, each client is assumed to have a small local data set, while the total number of clients is large, e.g., thousands or millions. In the cross-silo case, the total number of clients is small, for example, a dozen, but each client is assumed to have a larger local data set. In this paper, our running example is standard cross-silo differentially private FL (DPFL) where the clients communicate all updates to the server using SecAgg. [^1] Instead of considering any specific SecAgg implementation, in this work we mostly assume an idealised trusted aggregator. We discuss practical implementations in Appendix A.4. In this setting, the most useful DP protection granularity is typically something strictly more fine-grained than client-level: when clients are, e.g., different hospitals or banks, there are typically several individuals in a single clients’ local data set and the protection granularity needs to match the use case.
@@
While client-level granularity in DPFL is, at least in principle, straightforward to combine with SecAgg, more fine-grained granularities such as sample-level DP can present problems: using existing techniques one has to choose between i) having DDP guarantees with less noise due to SecAgg but with all clients using only a single local optimization step per FL round, and ii) having more noisy local DP (LDP) guarantees that do not formally benefit from SecAgg while allowing the clients to do more local optimization steps per FL round. Both of these options have significant drawbacks: the amount of server-client communications is typically one of the first bottlenecks that limit model training in FL, while LDP guarantees regularly require noise levels that heavily affect the resulting model utility. In this paper we show that this trade-off is not unavoidable but can be largely remedied by a simple new analysis of the problem.
@@
  • We present a novel and simple theoretical privacy analysis showing when we can increase the number of local optimization steps in FL using fine-grained DP granularity, while still benefiting from DDP guarantees using a trusted aggregator.
  • We demonstrate empirically that the proposed approach can lead to large utility benefits (in terms of prediction accuracy and loss) without requiring any changes to the underlying algorithms under both iid and heterogeneous client data splits. In our experiments with limited number of global communication rounds, using a convolutional neural network with Fashion MNIST data set as well as linear models on CIFAR-10 (using pre-trained model as feature extractor) and ACS Income data sets, we improve prediction accuracy roughly by $16\%$ points, $4\%$ points, and $4\%$ points, respectively (corresponding to $23\%$, $5\%$ and $5\%$ improvements).
  • Our results point to a mismatch between the current theoretical understanding of vanilla DPFL (standard DPFL with FedAvg) and practical results.
@@ 2 Related Work
There is a significant amount of existing work focusing on the general problem of combining DP with FL, although the focus has mostly been on the cross-device FL setting with user- or client-level DP. To the best of our knowledge, while the combination of DPFL with SecAgg is certainly not novel (see, e.g., Truex et al. 2019; Kairouz et al. 2019; Heikkilä et al. 2020; Stevens et al. 2022; Yang et al. 2023), there is very limited amount of work specifically on the privacy analysis when the clients do multiple local optimization steps with fine-grained DP and communicate the results via SecAgg. As far as we know, the only directly relevant analysis is by Noble et al. (2022), who consider the Gaussian mechanism for sample-level DP in the cross-device setting without secure aggregation. Our main theorem (Theorem 4.6) can be seen as extending their result on the Gaussian mechanism to other DP mechanisms that are more suitable for combining with actual secure aggregation protocols.  [^2] Note that (Truex et al., 2019, Algorithm 4) also state a similar specialised version for the Gaussian mechanism, i.e., they use several local optimization steps with sample-level DP and SecAgg in FL, while scaling the Gaussian noise for DDP considering all the clients. However, as also noted by Malekzadeh et al. (2021), the approach of Truex et al. (2019) would require a separate proof of privacy beyond what is actually provided in the paper.
Considering other existing work in more detail, we can distinguish some main lines of closely-related research. Besides the DDP approach we focus on, i.e., the combination of LDP noise with secure aggregation, the use of LDP mechanism by each client independently provides robust privacy with few assumptions, but generally results in worse utility (see, e.g., Kasiviswanathan et al. 2008; Truex et al. 2020). Another recently considered approach assumes a trusted server that is tasked with the DP noise addition, and does privacy accounting via the matrix mechanism (Zhang et al., 2024). This can be shown to improve the privacy-utility trade-off compared to standard DP-SGD under some assumptions, but is not directly compatible with the secure aggregation approach for DDP we focus on in this work.
@@
Looking at somewhat more closely-related papers in terms of the general setting, there is plenty of work proposing novel learning methods for FL, assuming sample-level DP and LDP noise with SecAgg for improved DDP guarantees. While the existing work only uses a single local optimization step (see, e.g., Heikkilä et al. 2020; Malekzadeh et al. 2021; Stevens et al. 2022; Yang et al. 2023), our analysis can be leveraged in this setting to enable running multiple local steps generally for many such methods without requiring any other changes to the algorithms.
@@
Another clear line of work has focused on introducing novel discrete DP mechanisms that can be used with additively homomorphic encryption techniques, which typically operate on the group of integers with modulo additions. Agarwal et al. (2018) proposed a binomial mechanism that provides DP using discrete binomial noise. Improving on the binomial mechanism, Canonne et al. (2020) proposed a discrete Gaussian mechanism, while Agarwal et al. (2021) introduced a Skellam mechanism and Chen et al. (2022b) a Poisson-binomial mechanism, both of which improve on the discrete Gaussian, e.g., by being infinitely divisible distributions: the sum of Skellam/Poisson-binomial distributed random variables is another Skellam/Poisson-binomial random variable. Our work is not focused on introducing new DP mechanisms, but our analysis allows for using many different DP noise mechanism. In particular, our analysis allows for using LDP noise with SecAgg including when using infinitely divisible DP mechanisms, such as the Skellam mechanism, with pseudo-gradients and fine-grained DP protection level.
While our main focus is on privacy accounting with SecAgg under limited communication budget, there has also been considerable effort by the community to reduce the amount of required communication further by applying quantization to the gradients (Agarwal et al., 2018; Kairouz et al., 2021; Agarwal et al., 2021; Chen et al., 2022b; Jin et al., 2020; Chaudhuri et al., 2022; Guo et al., 2023) or by compressing the updates sent by the clients (Triastcyn et al., 2021; Chen et al., 2022a). In principle, any such technique for compressing the model updates compatible with SecAgg can also be directly combined with our privacy analysis. In contrast, benefiting from gradient quantization is not entirely straightforward as in our case the model updates are pseudo-gradients and not gradients. We leave a detailed consideration and comparison of the possible methods for reducing the required communication budget beyond what is possible by pushing more optimization steps to the clients for future work.
@@
In summary, while many of the contributions cited above, e.g., novel DP mechanisms, are not limited to cross-device FL, all the experiments and use cases mentioned in the cited papers that are compatible with SecAgg and use multiple local steps only consider DDP with user- or client-level DP in cross-device FL. In contrast, we focus on more fine-grained DP granularities, namely on sample-level DP. As we discuss in Section 3, combining sample-level DP with multiple local steps and LDP noise with SecAgg for improved utility requires a novel privacy analysis. The main aim of this paper is to provide such an analysis.
While the currently existing theoretical convergence bounds for vanilla DPFL with FedAvg do not show any benefit from increasing the number of local steps in DPFL (see Malekmohammadi et al. 2024, Theorem 3.2), we empirically demonstrate the utility of our analysis in Section 5 after stating the results in Section 4. Our results clearly highlight the need for improving the theoretical analysis of standard DPFL over what is shown by Malekmohammadi et al. (2024) to understand when increasing the number of local steps is useful (compare this disagreement of empirical results and theory to the discussion by Mishchenko et al. 2022 on the provable usefulness of local steps in vanilla non-DP FL).
@@ 3 Background
Federated learning (FL, McMahan et al. 2017; Kairouz et al. 2019) is a collaborative learning setting, where the participants include a central server and clients holding some data. On each FL round, the server chooses a group of clients for an update and sends them the current model parameters. The chosen clients update their local model parameters by taking some amount of optimization steps using only their own local data, and then send an update back to the server. The server then aggregates the client-specific contributions to update the global model. We use the standard federated averaging update rule: assuming w.l.o.g. that clients $i=1,\dots,N$ have been selected at FL round $t$, and that client $i$ sends an update $\Delta^{(t)}_{i}$ (pseudo-gradient), the updated global model $\theta_{t}$ is given by
@@
3.1 Differential Privacy
We want to guarantee privacy of the trained model w.r.t. the training data, for which we use differential privacy (DP). Writing the space of possible data sets as $\mathcal{X}^{*}:=\cup_{n\in\mathbb{N}}\mathcal{X}^{n}$, we have the following:
@@ 3.1 Differential Privacy
(Dwork et al., 2006b; a) Let $\varepsilon>0$ and $\delta\in[0,1]$. A randomised algorithm ${\mathcal{A}}:\,\mathcal{X}^{*}\rightarrow\mathcal{O}$ is $(\varepsilon,\delta)$-DP if for every $x,x^{\prime}\in\mathcal{X}^{*}:x\simeq x^{\prime}$, and every measurable set $E\subset\mathcal{O}$,
@@
$H_{\alpha}(P\|Q):=\mathbb{E}_{t\sim Q}\left(\left[\frac{dP}{dQ}(t)-\alpha\right]_{+}\right)=\mathbb{E}_{t\sim Q}\left(\left[\frac{p(t)}{q(t)}-\alpha\right]_{+}\right),$ (2)
@@
It has been shown that a randomised algorithm ${\mathcal{A}}$ is $(\varepsilon,\delta)$-DP iff $\sup_{x\simeq x^{\prime}}H_{{\rm e}\hskip 1.0pt^{\varepsilon}}({\mathcal{A}}(x)\|{\mathcal{A}}(x^{\prime}))\leq\delta$ (Barthe et al., 2013; Barthe & Olmedo, 2013).
Our main results do not depend on the exact neighbourhood definition, but in all the experiments we use the add/remove relation or unbounded DP, that is, $x,x^{\prime}\in\mathcal{X}^{*}$ are neighbours, if $x$ can be transformed into $x^{\prime}$ by adding or removing a single protected unit from $\mathcal{X}$. For the protection granularity, although this is not a strict limitation, we focus on the common sample-level DP, i.e., a single protected unit corresponds to a single sample of data. As noted earlier, our analysis could be advantageous for anything more fine-grained than client-level DP, e.g., element-level DP (Asi et al., 2019) or even for individual-level when there are more than one individual in the clients’ local data.
@@
(Zhu et al., 2022) A pair of distributions $(P,Q)$ is a dominating pair for a stochastic algorithm ${\mathcal{A}}$, if for all $\alpha\geq 0$
@@
3.2 Problem with Local Steps in DPFL with SecAgg
While DP offers strict privacy protection, it comes at the cost of reduced model utility. This is especially true in the local DP (LDP) setting, where each client protects its own data independently of any other party (Kasiviswanathan et al., 2008). One well-known technique to improve model utility in DPFL has been to utilise secure aggregation (SecAgg) to turn LDP guarantees into distributed DP (DDP) guarantees that depend on multiple clients (see, e.g., Kairouz et al. 2019). [^3] Here, we distinguish between local DP (LDP, each client guarantees DP for it’s own contribution locally and in isolation), distributed DP (DDP, a group of clients jointly contribute to the resulting DP guarantees against outside adversaries without a trusted central party) and centralised DP (CDP, a trusted party guarantees DP for everyone). In effect, this means that when the clients are able to collaborate, they can jointly scale the locally used noise levels to reach a target DDP guarantee resulting in improved utility compared to the LDP guarantees. Even if the clients do not explicitly collaborate on calibrating the noise levels, the DDP guarantees resulting from using SecAgg will nevertheless be stronger compared to the LDP guarantees reached by each client individually.
However, naively combining fine-grained DP protection with SecAgg for DDP runs into problems, as we demonstrate in the rest of this section. For convenience, we give the pseudo-code for running standard DPFL with sample- and user-level DP protection in Appendix A.2.
Starting with the unproblematic case of client-level DP, writing $TA$ for an ideal trusted aggregator and using the well-known Gaussian mechanism (Dwork et al., 2006a) for simplicity, one can get DDP guarantees for any number of local optimization steps with the following update: